If a=-2 and b=16, then a^0 b^1/2 + a^-2 b^-1/2 is equal to...
A. -3/6 B. 1/32 C. 1/16 D. 65/16 E. 257/32

Answer by stanbon(75887)

If a=-2 and b=16, then a^0 b^1/2 + a^-2 b^-1/2 is equal to...
---
= 1*sqrt(16) + (1/4)(1/4)
= 4 + 1/16
= (16/16)+(1/16)
= 17/16
------
Cheers,
Stan H.
================
A. -3/6 B. 1/32 C. 1/16 D. 65/16 E. 257/32